# TsCA: On the Semantic Consistency Alignment via Conditional Transport for Compositional Zero-Shot Learning

### v1

@@ Introduction
Pairwise CT Distance.
Given the source distribution ${\bf P}_{1}$ and the target distribution ${\bf P}_{2}$, CT measures the distance by calculating the total transport costs bidirectionally, leading to the forward CT and backward CT, respectively. Denoting $c(\boldsymbol{x}_{n},\boldsymbol{y}_{m})\geq 0$ as a cost function to define the difference between points $\boldsymbol{x}_{n}$ and $\boldsymbol{y}_{m}$, the forward CT is measured as the expected transport cost of transporting all points from ${\bf P}_{1}$ to the target ${\bf P}_{2}$ set, and the backward CT reverses the transport direction. Mathematically, the CT distance between ${\bf P}_{1}$ and ${\bf P}2$ can be expressed as:
@@
$\mathrm{s.t.}\overrightarrow{{\bf T}}\mathds{1}^{N}=\boldsymbol{\theta},\overleftarrow{{\bf T}}\mathds{1}^{M}=\boldsymbol{\alpha},$
where we employ the cosine distance as the cost function, i.e., the closer the patch $\boldsymbol{x}_{n}$ and composition $\boldsymbol{y}_{m}$ are, the lower the transport cost. $\mathds{1}^{M}$ is the $M$ dimensional vector of ones. ${\bf T}$ is called the transport plan, which is to be learned to minimize the total distance. The forward transport plan, e.g., $\overrightarrow{t}_{nm}$ in $\overrightarrow{{\bf T}}$ describes the transport probability from the source point $\boldsymbol{x}_{n}$ to target point $\boldsymbol{y}_{m}$:
$\overrightarrow{t}_{nm}=\theta_{n}\frac{\alpha_{m}e^{s_{\psi}(\boldsymbol{x}_{n},\boldsymbol{y}_{m})}}{\sum_{m^{\prime}=1}^{M}\alpha_{m}{{}^{\prime}}e^{s_{\psi}(\boldsymbol{x}_{n},\boldsymbol{y}_{m^{\prime}})}},$ (9)
where $s_{\psi}$ denotes the similarity function with learnable parameter $\psi$, and we specify it as $s_{\psi}(\boldsymbol{x}_{n},\boldsymbol{y}_{m})=\frac{\boldsymbol{x}_{n}^{T}\boldsymbol{y}_{m}}{\text{exp}(\psi)}$. That is, the patch $\boldsymbol{x}_{n}$ will transport with higher probability if the composition $\boldsymbol{y}_{m}$ shares similar semantics with it. Similarly, we have the backward transport expressed as:
$\overleftarrow{t}_{mn}=\alpha_{m}\frac{\theta_{n}e^{s_{\psi}(\boldsymbol{y}_{m},\boldsymbol{x}_{n})}}{\sum_{n^{\prime}=1}^{N}\theta_{n^{\prime}}e^{s_{\psi}(\boldsymbol{y}_{m},\boldsymbol{x}_{n^{\prime}})}}.$ (10)
@@
Cycle Consistency Regularization.
@@
Due to the independent operation in Eq. 11, pairwise CTs may lead to potential inconsistencies and inefficiencies, as they fail to fully capture the global relationships among all sets involved. Cycle consistency, which has been extensively explored in domains such as image matching (Bernard et al. 2019), multi-graph alignment (Wang, Yan, and Yang 2021), and 3D pose estimation (Dong et al. 2019), offers a natural solution to this problem. By enforcing a closed-loop structure, cycle consistency can enhance the coherence of semantic relationships across all sets, ensuring more consistent and efficient alignment. Here, we denote the probability of the composition set being self-mapping back to its initial state as:
${{\bf T}}_{22}=\overrightarrow{{\bf T}}_{23}\cdot\overrightarrow{{\bf T}}_{31}\cdot\overrightarrow{{\bf T}}_{12},$ (12)
@@
where $\overrightarrow{{\bf T}}_{23}$ denotes the transport plan from source ${\bf P}_{2}$ to target ${\bf P}_{3}$. This guarantees each label is transferred with a probability of 1 during transportation among three sets. For a given input, the corresponding point from the composition set first interacts with key patches in the patch set, then moves to the relevant state and object points within the primitive set, and finally returns to the composition set. Accordingly, we derive the cycle-consistency constraint by:
$\mathcal{L}_{cyc}=\sum_{m=1}^{M}\boldsymbol{y}^{c}_{m}({\bf T}_{22}-\boldsymbol{I}),$ (13)
@@
where $\boldsymbol{y}^{c}_{m}$ denotes the one-hot binary label vector of the input image. For one thing, such a closed-loop transport constraint establishes a close link between independent pairwise CTs, which supports maintaining semantic consistency throughout the CT transport process. For another, it facilitates more robust composition representation learning, enhancing the model’s accuracy.
Primitive Decoupler.
@@
In light of the intricate entanglement between attributes and objects within an image, we devise a decoupling loss with the idea that visual representation $\boldsymbol{x}^{s}_{CLS}$ and $\boldsymbol{x}^{o}_{CLS}$ can be seen as state-expert and object-expert if their paired sub-concepts fail to be inferred:
$\mathcal{L}_{de}=\left\|cos(\boldsymbol{x}^{s}_{CLS},\boldsymbol{z}^{o}_{gt})\right\|+\left\|cos(\boldsymbol{x}^{o}_{CLS},\boldsymbol{z}^{s}_{gt})\right\|,$ (14)
@@
where $\boldsymbol{z}^{o}_{gt}$ and $\boldsymbol{z}^{s}_{gt}$ are textual representations for the ground-truth object class and attribute class, respectively. By mitigating the entanglement among visual representations, TsCA enhances its capacity to pinpoint correlative image representations that align with specific knowledge, thereby offering guidance on the textual sets prior and playing a complementary role in deriving a more accurate composition during inference.
Tranining and Inference
Training Objectives.
@@
Recalling that the probability weights $\boldsymbol{\alpha}$, $\boldsymbol{\beta}^{s}$, and $\boldsymbol{\beta}^{o}$, calculated during the construction of the three sets, capture the semantic similarity between the visual image and the corresponding textual features, this naturally defines the probability for predicting the labels of state $s$, object $o$, and composition $c$ for the image $x$:
$p(s|x)=\boldsymbol{\beta}^{s},\quad p(o|x)=\boldsymbol{\beta}^{o},\quad p(c|x)=\boldsymbol{\alpha}$ (15)
@@
Experiment
Experimental Settings
Datasets. The proposed TsCA is evaluated on three real-world CZSL benchmark datasets: MIT-states(Isola, Lim, and Adelson 2015), UT-Zappos(Yu and Grauman 2014) and C-GQA (Naeem et al. 2021). MIT-States comprises images of naturally occurring objects, with each object characterized by an accompanying adjective description. It comprises 53,753 images depicting 115 states and 245 objects. UT-Zappos is a fine-grained dataset consisting of different kinds of shoes with texture attributes, totaling 16 states and 12 objects. C-GQA, derived from the Stanford GQA dataset (Hudson and Manning 2019), features 453 states and 870 objects with over 9,500 compositions, making it the most pairs dataset for CZSL. We follow the split suggested by the previous work (Purushwalkam et al. 2019) to ensure fair comparisons and detailed statistics can be found in Tab. 1.
Table 1: Statistics of datasets
Dataset Train Validation Test
|Ys| |X| |Ys| |Yu| |X| |Ys| |Yu| |X|
MIT-States 1262 30338 300 300 10420 400 400 12995
UT-Zappos 83 22998 15 15 3214 18 18 2914
C-GQA 5592 26920 1252 1040 7280 888 923 5098
Table 2: CZSL comparisons in the close-world setting. The best results are in bold. The second best results are in blue.
@@
CLIP (Radford et al. 2021) 30.2 46.0 26.1 11.0 15.8 49.1 15.6 5.0 7.5 25.0 8.6 1.4
CoOp (Zhou et al. 2022a) 34.4 47.6 29.8 13.5 52.1 49.3 34.6 18.8 20.5 26.8 17.1 4.4
PromptCompVL (Xu et al. 2022) 48.5 47.2 35.3 18.3 64.4 64.0 46.1 32.2 —— —— —— ——
CSP (Nayak, Yu, and Bach 2023) 46.6 49.9 36.3 19.4 64.2 66.2 46.6 33.0 28.8 26.8 20.5 6.2
HPL (Wang et al. 2023a) 47.5 50.6 37.3 20.2 63.0 68.8 48.2 35.0 30.8 28.4 22.4 7.2
GIPCOL (Xu, Chai, and Kordjamshidi 2024) 48.5 49.6 36.6 19.9 65.0 68.5 48.8 36.2 31.9 28.4 22.5 7.1
DFSP (i2t) (Lu et al. 2023a) 47.4 52.4 37.2 20.7 64.2 66.4 45.1 32.1 35.6 29.3 24.3 8.7
DFSP (BiF) (Lu et al. 2023a) 47.1 52.8 37.7 20.8 63.3 69.2 47.1 33.5 36.5 32.0 26.2 9.9
DFSP (t2i) (Lu et al. 2023a) 46.9 52.0 37.3 20.6 66.7 71.7 47.2 36.0 38.2 32.0 27.1 10.5
PLID (Bao et al. 2023) 49.7 52.4 39.0 22.1 67.3 68.8 52.4 38.7 38.8 33.0 27.9 11.0
Troika (Huang et al. 2024) 49.0 53.0 39.3 22.1 66.8 73.8 54.6 41.7 41.0 35.7 29.4 12.4
CDS-CZSL (Li et al. 2024a) 50.3 52.9 39.2 22.4 63.9 74.8 52.7 39.5 38.3 34.2 28.1 11.1
Retrieval-Augmented (Jing et al. 2024) 50.0 53.3 39.2 22.5 69.4 72.8 56.5 44.5 45.6 36.0 32.0 14.4
TsCA 51.2 52.9 39.9 23.0 68.6 74.3 58.4 45.2 42.9 38.9 32.5 14.6
@@
CLIP (Radford et al. 2021) 30.1 14.3 12.8 3.0 15.7 20.6 11.2 2.2 7.5 4.6 4.0 0.3
CoOp (Zhou et al. 2022a) 34.6 9.3 12.3 2.8 52.1 31.5 28.9 13.2 21.0 4.6 5.5 0.7
PromptCompVL (Xu et al. 2022) 48.5 16.0 17.7 6.1 64.6 44.0 37.1 21.6 —— —— —— ——
CSP (Nayak, Yu, and Bach 2023) 46.3 15.7 17.4 5.7 64.1 44.1 38.9 22.7 28.7 5.2 6.9 1.2
HPL (Wang et al. 2023a) 46.4 18.9 19.8 6.9 63.4 48.1 40.2 24.6 30.1 5.8 7.5 1.4
GIPCOL (Xu, Chai, and Kordjamshidi 2024) 48.5 16.0 17.9 6.3 65.0 45.0 40.1 23.5 31.6 5.5 7.3 1.3
DFSP (i2t) (Lu et al. 2023a) 47.2 18.2 19.1 6.7 64.3 53.8 41.2 26.4 35.6 6.5 9.0 2.0
DFSP (BiF) (Lu et al. 2023a) 47.1 18.1 19.2 6.7 63.5 57.2 42.7 27.6 36.4 7.6 10.6 2.4
DFSP (t2i) (Lu et al. 2023a) 47.5 18.5 19.3 6.8 66.8 60.0 44.0 30.3 38.3 7.2 10.4 2.4
PLID (Bao et al. 2023) 49.1 18.7 20.0 7.3 67.6 55.5 46.6 30.8 39.1 7.5 10.6 2.5
Troika (Huang et al. 2024) 48.8 18.7 20.1 7.2 66.4 61.2 47.8 33.0 40.8 7.9 10.9 2.7
CDS-CZSL (Li et al. 2024a) 49.4 21.8 22.1 8.5 64.7 61.3 48.2 32.3 37.6 8.2 11.6 2.7
Retrieval-Augmented (Jing et al. 2024) 49.9 20.1 21.8 8.2 69.4 59.4 47.9 33.3 45.5 11.2 14.6 4.4
TsCA (w/o filter) 50.7 21.4 21.9 8.6 67.6 63.5 52.0 36.1 43.2 11.2 14.6 4.3
TsCA (w filter) 50.8 21.5 22.0 8.7 68.3 63.8 52.5 36.7 43.6 11.3 14.7 4.5
@@
Metrics. Following the common practice of prior works (Lu et al. 2023a), we utilize the standard evaluation protocols and assessed all results using four metrics in both closed-world and open-world scenarios. Concretely, S measures the best seen accuracy when calibration bias is +$\infty$+ ∞ and U denotes the accuracy specifically for unseen compositions when the bias is $\infty$- ∞. To provide an overall performance measure on both seen and unseen pairs, we also report the area under the seen-unseen accuracy curve (AUC) by varying the calibration bias from +$\infty$+ ∞ to $\infty$- ∞ and identify the point that achieves the best harmonic mean (H) between the seen and unseen accuracy. Among these, H and AUC are the core metrics for comprehensively evaluating the model.
@@
Implementation Details. The proposed TsCA and all baselines are implemented with a per-trained CLIP ViT-L/14 model in PyTorch (Paszke et al. 2019). For the UT-Zappos, the hyper-parameters $\lambda_{0}$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ in losses are set as 1, 0.1, 0.1, and 1. For the MIT-States, the hyper-parameters are set as 1, 0.01, 0.05, and 0.01. For the CGQA, the hyper-parameters are set as 1, 0.01, 0.03, and 0.01. During inference, $\gamma$ is set to 0.8 for UT-Zappos and 0.4 for MIT-States and C-GQA in the closed-world scenario and is adjusted to 0.4, 0.3, and 0.2 for each of the datasets in the open-world setting. The primitive adapters consist of two individual single-layer MLPs. All experiments are performed on a single NVIDIA RTX A6000 GPU. Please refer to the appendix for more details.
Comparision with State-of-the-Arts

### v3

@@ 3.3 Semantic Consistency Alignment
3.3.1 Pairwise CT Distance.
Given the source distribution ${\bf P}_{1}$ and the target distribution ${\bf P}_{2}$, CT measures the distance by calculating the total transport costs bidirectionally, leading to the forward CT and backward CT, respectively. Denoting $c(\bm{x}_{n},\bm{y}_{m})\geq 0$ as a cost function to define the difference between points $\bm{x}_{n}$ and $\bm{y}_{m}$, the forward CT is measured as the expected transport cost of transporting all points from ${\bf P}_{1}$ to the target ${\bf P}_{2}$ , and the backward CT reverses the transport direction. Mathematically, the CT distance between ${\bf P}_{1}$ and ${\bf P}2$ can be expressed as:
@@
$\mathrm{s.t.}\overrightarrow{{\bf T}}\mathds{1}^{N}=\bm{\theta},\overleftarrow{{\bf T}}\mathds{1}^{M}=\bm{\alpha},$
where we employ the cosine distance as the cost function, i.e., the closer the patch $\bm{x}_{n}$ and composition $\bm{y}_{m}$ are, the lower the transport cost. $\mathds{1}^{M}$ is the $M$ dimensional vector of ones. ${\bf T}$ is called the transport plan, which is to be learned to minimize the total distance. The forward transport plan, e.g., $\overrightarrow{t}_{nm}$ in $\overrightarrow{{\bf T}}$ describes the transport probability from the source point $\bm{x}_{n}$ to target point $\bm{y}_{m}$:
$\overrightarrow{t}_{nm}=\theta_{n}\frac{\alpha_{m}e^{s_{\psi}(\bm{x}_{n},\bm{y}_{m})}}{\sum_{m^{\prime}=1}^{M}\alpha_{m^{\prime}}e^{s_{\psi}(\bm{x}_{n},\bm{y}_{m^{\prime}})}},$ (9)
where $s_{\psi}$ denotes the similarity function with learnable parameter $\psi$, and we specify it as $s_{\psi}(\bm{x}_{n},\bm{y}_{m})=\frac{\bm{x}_{n}^{T}\bm{y}_{m}}{\text{exp}(\psi)}$. That is, the patch $\bm{x}_{n}$ will transport with higher probability if the composition $\bm{y}_{m}$ shares similar semantics with it. Similarly, we have the backward transport expressed as:
$\overleftarrow{t}_{mn}=\alpha_{m}\frac{\theta_{n}e^{s_{\psi}(\bm{y}_{m},\bm{x}_{n})}}{\sum_{n^{\prime}=1}^{N}\theta_{n^{\prime}}e^{s_{\psi}(\bm{y}_{m},\bm{x}_{n^{\prime}})}}.$ (10)
@@
3.3.2 Cycle Consistency Regularization.
@@
Due to the independent operation in Eq. 11, pairwise CTs may lead to potential inconsistencies and inefficiencies, as they fail to fully capture the global relationships among all sets involved. Cycle consistency, which has been extensively explored in domains such as image matching Bernard et al. (2019), multi-graph alignment Wang et al. (2021), and 3D pose estimation Dong et al. (2019), offers a natural solution to this problem. By enforcing a closed-loop structure, cycle consistency can enhance the coherence of semantic relationships across all sets, ensuring more consistent and efficient alignment. Here, we denote the probability of the composition set being self-mapping back to its initial state as:
${{\bf T}}_{22}=\overleftarrow{{\bf T}}_{21}\cdot\overrightarrow{{\bf T}}_{13}\cdot\overleftarrow{{\bf T}}_{32},$ (12)
@@
where $\overleftarrow{{\bf T}}_{21}$ from $\mathcal{CT}({\bf P}_{1},{\bf P}_{2})$ denotes the transport plan from ${\bf P}_{2}$ to ${\bf P}_{1}$. This guarantees each label is transferred with a probability of 1 during transportation among three sets. For a given input, the corresponding point from the composition set first interacts with key patches in the patch set, then moves to the relevant state and object points within the primitive set, and finally returns to the composition set. Accordingly, we derive the cycle-consistency constraint by:
$\mathcal{L}_{cyc}=\sum_{m=1}^{M}\bm{y}^{c}_{m}({\bf T}_{22}-\bm{I}),$ (13)
where $\bm{y}^{c}_{m}$ denotes the one-hot binary label vector of the input image.
@@
For one thing, such a closed-loop transport constraint establishes a close link between independent pairwise CTs, which supports maintaining semantic consistency throughout the CT transport process. For another, it facilitates more robust composition representation learning, enhancing the model’s accuracy.
3.3.3 Primitive Decoupler.
@@
In light of the intricate entanglement between attributes and objects within an image, we devise a decoupling loss with the idea that visual representation $\bm{x}^{s}_{CLS}$ and $\bm{x}^{o}_{CLS}$ can be seen as state-expert and object-expert if their paired sub-concepts fail to be inferred:
$\mathcal{L}_{de}=\left\|cos(\bm{x}^{s}_{CLS},\bm{z}^{o}_{gt})\right\|+\left\|cos(\bm{x}^{o}_{CLS},\bm{z}^{s}_{gt})\right\|,$ (14)
@@
where $\bm{z}^{o}_{gt}$ and $\bm{z}^{s}_{gt}$ are textual representations for the ground-truth object class and attribute class, respectively. By mitigating the entanglement among visual representations, TsCA enhances its capacity to pinpoint correlative image representations that align with specific knowledge, thereby offering guidance on the textual sets prior and playing a complementary role in deriving a more accurate composition during inference.
3.4 Tranining and Inference
3.4.1 Training Objectives.
@@ 3.4 Tranining and Inference
Recalling that the probability weights $\bm{\alpha}$, $\bm{\beta}^{s}$, and $\bm{\beta}^{o}$, calculated during the construction of the three sets, capture the semantic similarity between the visual image and the corresponding textual features, this naturally defines the probability for predicting the labels of state $s$, object $o$, and composition $c$ for the image $x$:
$p(s|x)=\bm{\beta}^{s},\quad p(o|x)=\bm{\beta}^{o},\quad p(c|x)=\bm{\alpha}$ (15)
@@
4 Experiment
4.1 Experimental Settings
Datasets. The proposed TsCA is evaluated on three real-world CZSL benchmark datasets: MIT-statesIsola et al. (2015), UT-ZapposYu and Grauman (2014) and C-GQA Naeem et al. (2021). MIT-States comprises images of naturally occurring objects, with each object characterized by an accompanying adjective description. It comprises 53,753 images depicting 115 states and 245 objects. UT-Zappos is a fine-grained dataset consisting of different kinds of shoes with texture attributes, totaling 16 states and 12 objects. C-GQA, derived from the Stanford GQA dataset Hudson and Manning (2019), features 453 states and 870 objects with over 9,500 compositions, making it the most pairs dataset for CZSL. We follow the split suggested by the previous work Purushwalkam et al. (2019) to ensure fair comparisons and detailed statistics can be found in Appendix.
@@ 4.1 Experimental Settings
CLIP Radford et al. (2021) 30.2 46.0 26.1 11.0 15.8 49.1 15.6 5.0 7.5 25.0 8.6 1.4
CoOp Zhou et al. (2022) 34.4 47.6 29.8 13.5 52.1 49.3 34.6 18.8 20.5 26.8 17.1 4.4
PromptCompVL Xu et al. (2022) 48.5 47.2 35.3 18.3 64.4 64.0 46.1 32.2 —— —— —— ——
CSP Nayak et al. (2023) 46.6 49.9 36.3 19.4 64.2 66.2 46.6 33.0 28.8 26.8 20.5 6.2
HPL Wang et al. (2023) 47.5 50.6 37.3 20.2 63.0 68.8 48.2 35.0 30.8 28.4 22.4 7.2
GIPCOL Xu et al. (2024) 48.5 49.6 36.6 19.9 65.0 68.5 48.8 36.2 31.9 28.4 22.5 7.1
DFSP (i2t) Lu et al. (2023) 47.4 52.4 37.2 20.7 64.2 66.4 45.1 32.1 35.6 29.3 24.3 8.7
DFSP (BiF) Lu et al. (2023) 47.1 52.8 37.7 20.8 63.3 69.2 47.1 33.5 36.5 32.0 26.2 9.9
DFSP (t2i) Lu et al. (2023) 46.9 52.0 37.3 20.6 66.7 71.7 47.2 36.0 38.2 32.0 27.1 10.5
PLID Bao et al. (2023) 49.7 52.4 39.0 22.1 67.3 68.8 52.4 38.7 38.8 33.0 27.9 11.0
Troika Huang et al. (2024) 49.0 53.0 39.3 22.1 66.8 73.8 54.6 41.7 41.0 35.7 29.4 12.4
CDS-CZSL Li et al. (2024) 50.3 52.9 39.2 22.4 63.9 74.8 52.7 39.5 38.3 34.2 28.1 11.1
Retrieval-Augmented Jing et al. (2024) 50.0 53.3 39.2 22.5 69.4 72.8 56.5 44.5 45.6 36.0 32.0 14.4
TsCA 51.2 52.9 39.9 23.0 68.7 75.8 58.5 46.1 43.8 38.9 33.1 15.2
Table 1: CZSL comparisons in the closed-world setting. The best results are in bold. The second best results are in blue.
@@
CLIP Radford et al. (2021) 30.1 14.3 12.8 3.0 15.7 20.6 11.2 2.2 7.5 4.6 4.0 0.3
CoOp Zhou et al. (2022) 34.6 9.3 12.3 2.8 52.1 31.5 28.9 13.2 21.0 4.6 5.5 0.7
PromptCompVL Xu et al. (2022) 48.5 16.0 17.7 6.1 64.6 44.0 37.1 21.6 —— —— —— ——
CSP Nayak et al. (2023) 46.3 15.7 17.4 5.7 64.1 44.1 38.9 22.7 28.7 5.2 6.9 1.2
HPL Wang et al. (2023) 46.4 18.9 19.8 6.9 63.4 48.1 40.2 24.6 30.1 5.8 7.5 1.4
GIPCOL Xu et al. (2024) 48.5 16.0 17.9 6.3 65.0 45.0 40.1 23.5 31.6 5.5 7.3 1.3
DFSP (i2t) Lu et al. (2023) 47.2 18.2 19.1 6.7 64.3 53.8 41.2 26.4 35.6 6.5 9.0 2.0
DFSP (BiF) Lu et al. (2023) 47.1 18.1 19.2 6.7 63.5 57.2 42.7 27.6 36.4 7.6 10.6 2.4
DFSP (t2i) Lu et al. (2023) 47.5 18.5 19.3 6.8 66.8 60.0 44.0 30.3 38.3 7.2 10.4 2.4
PLID Bao et al. (2023) 49.1 18.7 20.0 7.3 67.6 55.5 46.6 30.8 39.1 7.5 10.6 2.5
Troika Huang et al. (2024) 48.8 18.7 20.1 7.2 66.4 61.2 47.8 33.0 40.8 7.9 10.9 2.7
CDS-CZSL Li et al. (2024) 49.4 21.8 22.1 8.5 64.7 61.3 48.2 32.3 37.6 8.2 11.6 2.7
Retrieval-Augmented Jing et al. (2024) 49.9 20.1 21.8 8.2 69.4 59.4 47.9 33.3 45.5 11.2 14.6 4.4
TsCA (w/o filter) 50.7 21.5 22.0 8.6 69.7 63.3 52.0 37.0 43.7 11.3 14.6 4.3
TsCA (w filter) 50.8 21.7 22.3 8.7 69.8 63.4 52.2 37.1 44.3 11.4 14.7 4.5
@@
Metrics. Following the common practice of prior works Lu et al. (2023), we utilize the standard evaluation protocols and assessed all results using four metrics in both closed-world and open-world scenarios. Concretely, S measures the best seen accuracy when calibration bias is +$\infty$+ ∞ and U denotes the accuracy specifically for unseen compositions when the bias is $\infty$- ∞. To provide an overall performance measure on both seen and unseen pairs, we also report the area under the seen-unseen accuracy curve (AUC) by varying the calibration bias from +$\infty$+ ∞ to $\infty$- ∞ and identify the point that achieves the best harmonic mean (H) between the seen and unseen accuracy. Among these, H and AUC are the core metrics for comprehensively evaluating the model.
@@
Implementation Details. The proposed TsCA and all baselines are implemented with a per-trained CLIP ViT-L/14 model in PyTorch Paszke et al. (2019). For the UT-Zappos, the hyper-parameters $\lambda_{0}$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ in losses are set as 1, 0.1, 10, and 0.1. For the MIT-States, the hyper-parameters are set as 1, 0.01, 0.1, and 0.01. For the CGQA, the hyper-parameters are set as 1, 0.01, 0.3, and 0.01. During inference, $\gamma$ is set to 0.8 for UT-Zappos and 0.4 for MIT-States and C-GQA in the closed-world scenario and is adjusted to 0.4, 0.3, and 0.2 for each of the datasets in the open-world setting. The primitive adapters consist of two individual single-layer MLPs. All experiments are performed on a single NVIDIA RTX A6000 GPU. Please refer to the Appendix for more details.
4.2 Comparision with State-of-the-Arts